OpenStudy (anonymous):

Convert 300° into radians?

OpenStudy (rsadhvika):

\(\large 300 \times \frac{\pi}{180}\)

OpenStudy (anonymous):

5r/3

OpenStudy (rsadhvika):

u mean 5pi/3, then yes :)

OpenStudy (anonymous):

yeah lol

OpenStudy (rsadhvika):

lol ok i was worried u thought my pi as r ;p

OpenStudy (anonymous):

lol

OpenStudy (rsadhvika):

:D

OpenStudy (anonymous):

How about 5° into radians?

OpenStudy (rsadhvika):

do the same

OpenStudy (anonymous):

I'm getting 0.08

OpenStudy (rsadhvika):

multiply it with \(\frac{\pi}{180}\)

OpenStudy (rsadhvika):

try again

OpenStudy (rsadhvika):

\(\large 5 \times \frac{\pi}{180}\)

OpenStudy (anonymous):

is it π\15

OpenStudy (rsadhvika):

\(\large 5 \times \frac{\pi}{180}\) \(\large \cancel{5} \times \frac{\pi}{\cancel{180}^{36}}\) \(\large \frac{\pi}{36}\) radians

OpenStudy (anonymous):

omg! , Let me see for an example if I to Convert π\10 radians into degrees, I do the same thing?

OpenStudy (rsadhvika):

radiance to degrees, u multiply it wid \(\large \frac{180}{\pi}\)

OpenStudy (anonymous):

Which will bring up to 10 degrees?

OpenStudy (anonymous):

i mean not 10, 18

OpenStudy (rsadhvika):

Correct :) you got it !!

OpenStudy (rsadhvika):

pi/10 radians = 18 degrees

OpenStudy (anonymous):

Awwwwww yeah!!

OpenStudy (rsadhvika):

good job ! brb need to run an errand cya

OpenStudy (anonymous):

cya!